OpenStudy (anonymous):

if m+n=7 and 2n-3m=6 what is the value of the 3n-2m

jimthompson5910 (jim_thompson5910):

m+n=7 is the same as n+m = 7

jimthompson5910 (jim_thompson5910):

now add the equations and tell me what you get n+m = 7 2n-3m=6 ---------- ???

OpenStudy (anonymous):

i got m=8/5

jimthompson5910 (jim_thompson5910):

you don't need to solve for m or n

jimthompson5910 (jim_thompson5910):

you just need to find 3n - 2m

OpenStudy (anonymous):

how

jimthompson5910 (jim_thompson5910):

add the two equations

OpenStudy (anonymous):

omg

OpenStudy (anonymous):

i got it

jimthompson5910 (jim_thompson5910):

glad it's clicking now